OpenStudy (anonymous):

need help help me out Find 3 consecutive even numbers where the product of the smaller two numbers is 64 less than the square of the largest number.

OpenStudy (anonymous):

x+x+2+x+4=(x)*(x+2)=(x+4)^2-64 I think